OpenStudy (anonymous):

can someone please help get the answer Solve 2^(x + 3) = 7 Answer a.x ≈ –5.19 b.x ≈ 2.81 c.x ≈ –0.19 d.x ≈ 0.9

OpenStudy (anonymous):

You can simplify it better by: 2^x + 2^3 = 7 Or go to http://www.wolframalpha.com/input/?i=solve+2^%28x%2B3%29%3D7 The answer is C.

OpenStudy (anonymous):

I havent done "log" in a looooong time, so sorry i cant help with that part

OpenStudy (apoorvk):

@sam that should be "2^x *2^3" and not '+'.

OpenStudy (anonymous):

right, sorry

OpenStudy (apoorvk):

np

OpenStudy (anonymous):

\[2^{x} \times 2^{3} = 7 \]

OpenStudy (anonymous):

I GOT IT!! Ok, so: \[2^{x} \times 8 =7\] \[2^{x}= 7/8\] \[x=\log(7/8) \div \log(2) ; x=-0.19\]